OpenStudy (anonymous):

What is the equation of a line that passes through (-6, 2) and has a slope of -1/2?

OpenStudy (zpupster):

you will need point slope form: \[y-y1=m(x-x1)\] we know our slope=m=-1/2 and x1 = -6 y1 = 2 plug those in

OpenStudy (anonymous):

y-2=-1/2(x-6)

OpenStudy (anonymous):

Is that the answer?

OpenStudy (zpupster):

not yet we should now put it in slope intercept form \[y=mx+b\] can you do that

OpenStudy (anonymous):

So... is it -1/2x-1

OpenStudy (zpupster):

also there was a mistake with the sign y-2=-1/2(x-6) should be y-2=-1/2(x-(-6)) y-2 = -1/2(x+6)

OpenStudy (anonymous):

so its y=-1/2x+5?

OpenStudy (zpupster):

y-2 = -1/2(x+6) distribute the -1/2 y-2 = -1/2x -3 add 2 both sides y=-1/2x-1

OpenStudy (zpupster):

here you see our point on the line

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what would the equation be?

OpenStudy (zpupster):

well you can see it above or you can see it on the graph

OpenStudy (anonymous):

i have four choices a:y=x/2-2 b:y=-x/2-1 c:y=2x+1 d:y=-2x-1